If money is worth 6% compounded bi-monthly (6), find the present value of R 125 700 for 2 years?

1
Expert's answer
2021-11-25T09:28:13-0500

compounded bi-monthly , n =6


"PV = \frac{FV}{(1+i)^n}"


"=\frac{R125,700}{(1+0.06)^{6×2}}"


"=\frac{125,700}{(1.06)^{12}}"


"=R62469.05"